OpenStudy (anonymous):

T=2 pie square root L/G L(feet) .5 1.0 1.5 T(sec.) .78 1.11 1.36 solve for g

OpenStudy (anonymous):

just plug the numbers in, the first one is g=2(pi/.78)^2

OpenStudy (anonymous):

if you mulitpy that it comes out to 2.4492 so do u round after that

OpenStudy (anonymous):

Where did the ^2 come into play? I would solve fo G and get g = 2 pi sqrt(L)/T So g = 2 * pi * sqrt(L) / T